OpenStudy (anonymous):

Consider the line y=−5/3x+1 . Find the equation of the line that is parallel to this line and passes through the point 6, 3 . Find the equation of the line that is perpendicular to this line and passes through the point 6, 3

OpenStudy (melodycrazysouthern):

um ull haft to look it up caz theres a million ways

OpenStudy (alexandervonhumboldt2):

A: parallel lines have same slopes. the slope of given line is -5/3. thus the slope of line parallel to it is also -5/3. now we have y=-5/3x+b. substitute the given coordinate: 3=-5/3*6 + b b=3+10 b=13 thus the equation is y=-5/3x+13

OpenStudy (alexandervonhumboldt2):

B: perpendicular lines have negative reciprocal slopes. the slope of given line is -5/3. thus the slope of line perpendicular to it is 3/5. now we have y=3/5x+b substitute given point: 3=3/5*6+b b=-0.6 thus the equation is y=3/5x-0.6